A merchant wishes to blend 200 pounds of coffee worth $3.20 a pound from two mixtures: one at $3.00 a pound and the other at $3.25 a pound. How many pounds of each mixture should he use?
Answer by TimothyLamb(4379)

x = pounds of $3.00 mixture
y = pounds of $3.25 mixture
---
x + y = 200
300x + 325y = 320*200
---
put the system of linear equations into standard form
---
x + y = 200
300x + 325y = 64000
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = pounds of $3.00 mixture = 40
y = pounds of $3.25 mixture = 160
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---